• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 01:40:57 on June18,2010

    Entry number 327158

    keyword=Parity feedback and beam recovery

    Parity feedback moniotor was not working, bob reduced the gains for 70 uA and PANGUIN was cutting the data for lobeam. Luis called and helpd to parity feedback working. Its working now. We have started teh beam recovery procedure.